OpenStudy (zenmo):

Determine whether the sequence converges or diverges. If it converges, find the limit.

OpenStudy (zenmo):

\[\left\{ \frac{ (2n-1)! }{ (2n+1)! } \right\}\]

OpenStudy (zarkon):

\[(2n+1)!=(2n+1)(2n)(2n-1)!\]

OpenStudy (zenmo):

Why are we multiplying the denominator with (2n)? Do we look at the 1st term, such as (2n + 1) and that is where we get the (2n) from?

ganeshie8 (ganeshie8):

Notice that 5! = 5*(5-1)!

ganeshie8 (ganeshie8):

In general : n! = n*(n-1)!